### v1

@@ 5.2 Localization results
Table 2: Visual localization results on the MGL dataset. $\uparrow$ means higher is better, $\downarrow$ means lower is better. The localization results of "Retrieval" and "Refinement" are taken from OrienterNet (Sarlin et al., 2023). We use bold font to indicate the best results and underline to indicate the second-best results.
@@
OrienterNet 10.69 42.12 54.01 19.22 48.56 63.59 12.00 28.71 5.5
OSMLoc-S 14.04 44.00 57.31 21.84 53.38 68.41 11.09 26.09 6.2
OSMLoc-B 15.30 46.05 59.40 21.79 56.57 71.29 10.04 22.91 5.0
MGL dataset: Table. 2 shows the single-image visual localization results on the MGL dataset. Notably, our OSMLoc significantly outperforms all baseline methods by a large margin across all position and orientation recall thresholds. Specifically, OSMLoc-S reduces the APE and the AOE by approximately $0.91\text{m}/2.62^{\circ}$ and improves the $5\text{m}/5^{\circ}$ recall by $3.30\%/4.82\%$ over the SOTA baseline method, OrienterNet (Sarlin et al., 2023); OSMLoc-B achieves the best performance, with a $1.96\text{m}/5.80^{\circ}$ reduction in APE and AOE and a $5.39\%/7.70\%$ improvement in the $5\text{m}/5^{\circ}$ recall. With comparable inference speed, the improvement mainly results from incorporating geometric and semantic guidance into the I2O localization framework rather than enlarging the model. Furthermore, Fig. 7 shows the recall curves for top candidates at different thresholds, where OSMLoc-B achieves the highest recall across various settings. These quantitative results demonstrate that our method significantly enhances prediction accuracy and robustness, as evidenced by consistent gains in position and orientation recall across different thresholds and top candidates.
@@
KITTI dataset: Extensive experiments are conducted on the KITTI dataset to demonstrate the generalizability of our method in self-driving scenarios. The quantitative results are shown in Table. 4. Without fine-tuning, our OSMLoc outperforms baseline methods trained on the KITTI dataset by a large margin, demonstrating its strong generalization capability and robustness. Furthermore, compared with the methods (Shi and Li, 2022; Sarlin et al., 2023) trained on the MGL dataset, our OSMLoc achieves much higher position and orientation recall. These quantitative results highlight that our method excels in accuracy, robustness, and generalization, making it suitable for versatile downstream tasks. Fig. 9 deploys additional qualitative results, where our method produces cleaner likelihood maps and more accurate camera pose predictions.
Table 4: Quantitative results on the KITTI dataset. We use bold font to indicate the best results and underline to indicate the second-best results..
@@
OSMLoc-S 65.75 93.79 96.74 24.50 59.53 72.67 33.41 75.67 91.67
OSMLoc-B 66.71 94.26 97.04 27.14 62.09 73.80 37.43 79.69 92.61
@@
5.3 Ablation studies and Analysis
Firstly, we conduct ablation studies to investigate the influence of the foundational model-based image encoder. As shown in Table. 5, the foundation model boosts the localization performance with its rich prior knowledge. Additionally, utilizing the larger DINOv2-B leads to continuous improvement in localization accuracy, but it also increases the computational complexity and reduces the efficiency. Table. 6 shows the ablation study results on the depth guidance and semantic guidance. In variant $[A]$, the framework is supervised by the pose label $\hat{\boldsymbol{p}}$ only. Variant $[B]$ utilizes the geometric guidance from the Depth Anything and introduces the depth loss $\mathcal{L}_{depth}$ into the framework. The localization results show that geometric guidance is beneficial for accurate localization. Variant $[C]$ utilizes the semantic embeddings $\mathrm{M}_{sem}$ from the OSM data and introduces the semantic alignment loss $\mathcal{L}_{sem}$, which also boosts the PR performance. Finally, variant $[D]$ demonstrates that jointly learning the pose, geometric and semantic achieves the best localization performance.
Table 5: Ablation studies on the foundation model of image encoder.
@@ 5.3 Ablation studies and Analysis
ResNet-101 12.90 43.27 56.05 20.48 52.28 66.42
DINOv2-S 14.04 44.00 57.31 $\boldsymbol{21.84}$ 53.38 68.41
DINOv2-B $\boldsymbol{15.30}$ $\boldsymbol{46.05}$ $\boldsymbol{59.40}$ 21.79 $\boldsymbol{56.57}$ $\boldsymbol{71.29}$
Table 6: Ablation studies on depth guidance and semantic guidance. OSMLoc-S is selected as the baseline. The $\mathcal{L}_{pose}$ means the pose classification loss. $\mathcal{L}_{depth}$ means the depth distribution loss. The $\mathcal{L}_{sem}$ means the semantic alignment loss.
@@
Failure cases: The localization accuracy of our OSMLoc heavily depends on the quality, quantity and distinctiveness of the surrounding objects in the image. When the camera’s FoV is obstructed by skyscrapers or the surrounding environment is monotonous, our method may struggle to accurately estimate the pose. Fig. 10 shows some failure cases under different conditions. In the first and second cases, useful information is limited due to the buildings obstructing the view, which makes it difficult for the model to distinguish the exact pose along the roadside. In the third and fourth cases, the model fails to localize itself with a single image, as the objects in the vehicle’s forward direction are extremely similar. Sequential localization with the motion pattern can greatly improve the accuracy, as discussed in Sec. 5.4.
@@ 5.4 Application in sequential localization
Following previous approaches (Chen et al., 2021; Zhou et al., 2021), we utilize the Monte Carlo Localization (MCL) framework with the particle filter to solve the sequential localization problem. For the first frame, we initialize $N$ particles onto the top-$N$ pose candidates and each particle represents a pose hypothesis $\boldsymbol{p}_{1}^{i}=(x_{1}^{i},y_{1}^{i},\theta_{1}^{i})$. When the camera moves, the pose of each particle is updated based on the motion model with the control input $\mathbf{u}$. The expected observation from the predicted pose of each particle is then compared with the actual observation acquired by the camera to update the weight of each sample. Particles are resampled based on the weight distribution and reduced for efficiency when the effective sample number is less than a predefined threshold. After several iterations, the particles gradually converge around the GT pose.
@@
[…]
where $\eta$ is the normalized constant, $p(\boldsymbol{p}_{t}|\boldsymbol{u}_{t},\boldsymbol{p}_{t-1})$ is the motion model, and the $p(\boldsymbol{S}_{t}|\boldsymbol{p}_{t},\mathcal{M})$ is the observation model. We use the standard motion model in (Thrun, 2002) and focus on the observation model. For each particle $i$, we query the matching score $\boldsymbol{S}_{t}(x_{t}^{i},y_{t}^{i},\theta_{t}^{i})$ from the score volume and approximate the likelihood $p(\boldsymbol{S}_{t}|\boldsymbol{p}_{t}^{i},\mathcal{M})$ as a Gaussian observation model:
@@
$\hat{\boldsymbol{p}}_{t}=\frac{\sum_{i=1}^{N}w_{t}^{i}\boldsymbol{p}_{t}^{i}}{\sum_{i=1}^{N}w_{t}^{i}}.$ (16)
During implementation, the number of particles is set to $N=1000$ at initialization and reduced to 200 if converged. The sensitivity parameter $\varsigma$ is set to 2.0. The minimal and maximal sequence lengths are set to 3 and 10 respectively. To simulate the real motion pattern, random Gaussian noise $\boldsymbol{\epsilon}_{p}$ is added to the GT pose. For the MGL dataset and CC validation benchmark, we sub-sample the sequences and ensure that frames are spaced by at least 4m, and $\epsilon_{x},\epsilon_{y}\sim\mathcal{N}(0,0.5\text{m}),\epsilon_{\theta}\sim\mathcal{N}(0,0.1\pi)$. For the KITTI dataset, we use the raw sequences and $\epsilon_{x},\epsilon_{y}\sim\mathcal{N}(0,0.05\text{m}),\epsilon_{\theta}\sim\mathcal{N}(0,0.01\pi)$.
@@
Table 7: Sequential localization results on the MGL dataset and CC validation benchmark. We use bold font to indicate the best results. “Detroit", “Munich", “Taipei" and “Brisbane" mean the city subsets of the CC validation benchmark. “single" means the single-image localization and “seq" means the sequential localization. "Ori" is short for OrienterNet.
@@
Table 8: Sequential localization results on the KITTI dataset. We use bold font to indicate the best results. “single" means the single-image localization and “seq" means the sequential localization.
@@ 6 Conclusion
In this paper, we present the OSMLoc, a brain-inspired single image-to-OpenStreetMap (I2O) visual localization framework. The proposed method integrates the recent foundation model Depth Anything for geometric guidance, and fully exploits semantic information from the OSM data to enhance the perception capability of the image model. A novel cross-area and cross-condition (CC) validation benchmark is proposed to evaluate the accuracy, robustness and generalizability of I2O visual localization methods. Experiments on the MGL dataset and CC validation benchmark have demonstrated that our OSMLoc achieves higher localization accuracy, while extensive experiments on the KITTI dataset highlight its application value in the autonomous driving area. Sequential localization experiments underscore the potential practical value. We hope that our OSMLoc will benefit the relevant communities in the field.
7 Acknowledgment
This study is supported by the National Natural Science Foundation Project (No. 42201477, No. 42130105).

### v3

@@ 5.2 Localization results
Table 2: Visual localization results on the MGL dataset. $\uparrow$ means higher is better, $\downarrow$ means lower is better. The localization results of "Retrieval" [65] and "Refinement" [64] are taken from OrienterNet [21]. We use bold font to indicate the best results and underline to indicate the second-best results.
@@
OrienterNet 10.69 42.12 54.01 19.22 48.56 63.59 12.00 28.71 10.77
AnyLoc 0.06 0.31 0.78 0.52 1.78 2.93 55.32 87.78 13.10
SNAP 9.11 37.51 49.08 19.49 48.04 62.23 13.20 31.09 12.18
OSMLoc-S 14.04 44.00 57.31 21.84 53.38 68.41 11.09 26.09 11.13
OSMLoc-B 15.30 46.05 59.40 21.79 56.57 71.29 10.04 22.91 9.27
MGL dataset: Table. 2 shows the single-image visual localization results on the MGL dataset. Notably, our OSMLoc significantly outperforms all baseline methods by a large margin across all position and orientation recall thresholds. Specifically, OSMLoc-S reduces the APE and the AOE by approximately $0.91\text{m}/2.62^{\circ}$ and improves the $5\text{m}/5^{\circ}$ recall by $3.30\%/4.82\%$ over the SOTA baseline method, OrienterNet [21]; OSMLoc-B achieves the best performance, with a $1.96\text{m}/5.80^{\circ}$ reduction in APE and AOE and a $5.39\%/7.70\%$ improvement in the $5\text{m}/5^{\circ}$ recall. With comparable inference speed, the improvement mainly results from incorporating geometric and semantic guidance into the I2O localization framework rather than enlarging the model. Furthermore, Fig. 7 shows the recall curves for top candidates at different thresholds, where OSMLoc-B achieves the highest recall across various settings. These quantitative results demonstrate that our method significantly enhances prediction accuracy and robustness, as evidenced by consistent gains in position and orientation recall across different thresholds and top candidates.
@@
KITTI dataset: Extensive experiments are conducted on the KITTI dataset to demonstrate the generalizability of our method in self-driving scenarios. The quantitative results are shown in Table. 4. Without fine-tuning, our OSMLoc outperforms baseline methods trained on the KITTI dataset by a large margin, demonstrating its strong generalization capability and robustness. Furthermore, compared with the methods [64, 65, 32, 54] trained on the MGL dataset, our OSMLoc achieves much higher position and orientation recall. These quantitative results highlight that our method excels in accuracy, robustness, and generalization, making it suitable for versatile downstream tasks. Moreover, our OSMLoc-S reduces the absolute latitudinal error (ALatE), longitudinal error (ALonE) and orientation error (AOE) from 1.35m/6.64m/2.50 to 1.28m/5.70m/1.99, respectively. Meanwhile, the OSMLoc-B achieves the lowest localization error at 1.15m/5.42m/1.99. Fig. 9 deploys additional qualitative results, where our method produces cleaner likelihood maps and more accurate camera pose predictions.
Table 4: Quantitative results on the KITTI dataset. We use bold font to indicate the best results and underline to indicate the second-best.
@@
AnyLoc 3.12 9.57 15.73 1.38 4.49 7.78 5.46 14.76 24.64
SNAP 41.05 84.02 92.64 22.86 53.53 64.56 25.32 60.63 79.67
OSMLoc-S 65.75 93.79 96.74 24.50 59.53 72.67 37.43 79.69 91.67
OSMLoc-B 66.60 95.05 97.41 29.40 63.86 74.52 35.57 78.28 92.96
@@
Table 5: Ablation studies on the foundation model of image encoder. The "FLOPs" metric includes only the computational complexity of the image encoder for fair comparison. ResNet101-FPN*: Image encoder employed by OrienterNet.
@@
ResNet-101 12.90 43.27 56.05 20.48 52.28 66.42 52.21G
DINOv2-S 14.04 44.00 57.31 $\boldsymbol{21.84}$ 53.38 68.41 $\boldsymbol{48.16}$G
DINOv2-B $\boldsymbol{15.30}$ $\boldsymbol{46.05}$ $\boldsymbol{59.40}$ 21.79 $\boldsymbol{56.57}$ $\boldsymbol{71.29}$ 154.10G
Table 6: Ablation studies on geometric guidance and semantic guidance. OSMLoc-S is selected as the baseline. "Geo" refers to the geometric guidance for camera-to-BEV transformation, while "Sem" denotes the semantic guidance from the OSM embeddings for cross-modal alignment.
@@
Table 7: Computational complexity, parameters and inference speed.
Method #Params FLOPS FPS
OrienterNet 54.94M 89.63G 10.77
SNAP 35.95M 70.17G 12.18
AnyLoc 35.38M 71.21G 13.10
OSMLoc-S 37.57M 85.57G 11.13
OSMLoc-B 110.60M 190.68G 9.27
5.3 Ablation studies and Analysis
Impact of the foundation model: Firstly, we conduct ablation studies to investigate the influence of the foundational model-based image encoder, as shown in Table. 5. For the ResNet101-FPN as in OrienterNet [21], we pre-build the pseudo normalized disparity ground-truth $\hat{\boldsymbol{t}}$ with the DAM-B model and utilize it as geometric guidance. Semantic guidance is obtained from OSM semantic embeddings, as in our approach. With the geometric and semantic guidance, the ResNet101-FPN based variant achieves better localization accuracy than the naive one in the OrienterNet [21], which demonstrates the application value of geometric and semantic guidance on general I2O localization frameworks. Furthermore, replacing the original ResNet101-FPN model with the DINOv2-S encoder boosts the localization accuracy with even less computational complexity, which deploys the powerful prior knowledge from foundational models to the I2O localization task. Utilizing the larger DINOv2-B leads to continuous improvement in localization accuracy, but it also increases the computational complexity and reduces the efficiency.
[…]
Impact of loss weights: We investigate the effect of the hyperparameters $\lambda_{1},\lambda_{2},\lambda_{3}$, which weight the pose loss $\mathcal{L}_{pose}$, disparity loss $\mathcal{L}_{dis}$, and semantic loss $\mathcal{L}_{sem}$, by experimenting with different configurations, as reported in Table. 8. Variant $[A]$ shows that the vanilla framework with $\mathcal{L}_{pose}$ only, while variants $[B]$ and $[C]$ with disparity loss $\mathcal{L}_{dis}$ and semantic loss $\mathcal{L}_{sem}$ achieve consistent improvement on localization accuracy. As shown in variants $[D],[E],[F]$, joint loss of pose, disparity and semantic loss boosts the performance, and set $(\lambda_{1},\lambda_{2},\lambda_{3})=(1,20,10)$ achieves the best result.
Table 8: Ablation studies on the loss weights. OSMLoc-S serves as the baseline, where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ denote the weights for the pose loss $\mathcal{L}_{pose}$, disparity loss $\mathcal{L}_{dis}$, and semantic loss $\mathcal{L}_{sem}$, respectively.
[…]
@1m @3m @5m @1 @3 @5
$[A]$ 1.0 0.0 0.0 11.52 42.54 55.53 20.80 51.44 65.79
$[B]$ 1.0 1.0 0.0 12.33 43.22 55.15 21.22 51.74 66.01
$[C]$ 1.0 0.0 1.0 13.20 43.43 56.16 20.43 51.08 66.16
$[D]$ 1.0 1.0 1.0 12.36 43.95 56.94 20.85 51.13 66.68
$[E]$ 10.0 10.0 10.0 12.89 44.05 55.94 21.00 53.87 67.84
$[F]$ 1.0 20.0 10.0 14.04 44.00 57.31 21.84 53.38 68.41
Computational complexity: We evaluate computational complexity, parameter count, and inference speed using image inputs of $512\times 512$ and map inputs of $256\times 256$. As shown in Table. 7, despite incorporating the DINOv2 encoder, our OSMLoc-S model has fewer parameters, lower computational complexity, and slightly faster inference speed compared to OrienterNet. Even the larger OSMLoc-B model maintains real-time performance with nearly 10 FPS, which suggests that the performance improvement stems from the strong prior knowledge embedded in the foundation model rather than from a heavier or more complex architecture. To enable lightweight online deployment on edge devices, an alternative is to adopt efficient vision backbones such as MobileSAM [69] for extracting semantic- and geometry-aware features. Moreover, knowledge distillation [70] also offers a promising direction to transfer performance from heavy teacher models to compact student networks.
@@ 5.3 Ablation studies and Analysis
Failure cases: The localization accuracy of our OSMLoc heavily depends on the quality, quantity and distinctiveness of the surrounding objects in the image. When visual information is limited, for example, when the camera’s field of view is obstructed by skyscrapers or dominated by monotonous surroundings, our method may struggle to estimate the pose accurately. Fig. 10 shows some failure cases under different conditions. In the first and second cases, the cameras are heavily occluded by tall nearby buildings, resulting in images that contain little more than the road and the facade of a single structure. Due to the repetitive visual patterns along the building-lined roadside, where similar scenes appear at multiple locations, the model cannot confidently pinpoint the query’s exact position, resulting in high localization ambiguity in the entire region. The third row, captured in an autonomous driving setting, presents a straight street bordered by visually repetitive buildings on both sides, a typical scene in many urban and suburban environments. This structural redundancy introduces significant uncertainty in determining the precise location. The last row shows a highway scene with few structured features or distinctive objects, resulting in high uncertainty along the vegetation-lined path.
@@ 5.4 Application in sequential localization
Following previous approaches [71, 20], we utilize the Monte Carlo Localization (MCL) framework with the particle filter to solve the sequential localization problem. For the first frame, we initialize $N_{p}$ particles onto the top-$N_{p}$ pose candidates and each particle represents a pose hypothesis $\boldsymbol{p}_{1}^{i}=(x_{1}^{i},y_{1}^{i},\theta_{1}^{i})$. When the camera moves, the pose of each particle is updated based on the motion model with the control input $\mathbf{u}$. The expected observation from the predicted pose of each particle is compared with the actual observation acquired by the camera to update the weight of each sample. Particles are resampled based on the weight distribution and reduced for efficiency when the effective sample number falls below a predefined threshold. After several iterations, the particles gradually converge around the GT pose.
@@
$\int p(\boldsymbol{p}_{t}|\boldsymbol{u}_{t},\boldsymbol{p}_{t-1})p(\boldsymbol{p}_{t-1}|\boldsymbol{S}_{1:t-1},\boldsymbol{u}_{1:t-1})\,d\boldsymbol{p}_{t-1},$ (6)
where $\eta$ is the normalized constant, $p(\boldsymbol{p}_{t}|\boldsymbol{u}_{t},\boldsymbol{p}_{t-1})$ is the motion model, and the $p(\boldsymbol{S}_{t}|\boldsymbol{p}_{t},\mathcal{M})$ is the observation model. We use the standard motion model in [72] and focus on the observation model. For each particle $i$, we query the matching score $\boldsymbol{S}_{t}(x_{t}^{i},y_{t}^{i},\theta_{t}^{i})$ from the score volume and approximate the likelihood $p(\boldsymbol{S}_{t}|\boldsymbol{p}_{t}^{i},\mathcal{M})$ as a Gaussian observation model:
@@
$\hat{\boldsymbol{p}}_{t}=\frac{\sum_{i=1}^{N_{p}}w_{t}^{i}\boldsymbol{p}_{t}^{i}}{\sum_{i=1}^{N_{p}}w_{t}^{i}}.$ (6)
During implementation, the number of particles is set to $N_{p}=1000$ at initialization and reduced to 200 if converged. The sensitivity parameter $\varsigma$ is set to 2.0. The minimal and maximal sequence lengths are set to 3 and 10 respectively. To simulate the real motion pattern, random Gaussian noise $\boldsymbol{\epsilon}_{p}$ is added to the GT pose. For the MGL dataset and CC validation benchmark, we sub-sample the sequences and ensure that frames are spaced by at least 4m, and $\epsilon_{x},\epsilon_{y}\sim\mathcal{N}(0,0.5\text{m}),\epsilon_{\theta}\sim\mathcal{N}(0,0.1\pi)$. For the KITTI dataset, we use the raw sequences and $\epsilon_{x},\epsilon_{y}\sim\mathcal{N}(0,0.05\text{m}),\epsilon_{\theta}\sim\mathcal{N}(0,0.01\pi)$.
@@
Table 9: Sequential localization results on the MGL dataset and CC validation benchmark. We use bold font to indicate the best results. "Detroit", "Munich", "Taipei" and "Brisbane" mean the city subsets of the CC validation benchmark. "single" means the single-image localization and "seq" means the sequential localization. "Ori" is short for OrienterNet.
@@
Table 10: Sequential localization results on the KITTI dataset. We use bold font to indicate the best results. "single" means the single-image localization and "seq" means the sequential localization.
@@ 6 Conclusion
In this paper, we present the OSMLoc, a brain-inspired single image-to-OpenStreetMap (I2O) visual localization framework. The proposed method fuses the recent foundation model Depth Anything for geometric guidance, and fully exploits semantic information from the OSM data to enhance the perception capability of the image model. A novel cross-area and cross-condition (CC) validation benchmark is proposed to evaluate the accuracy, robustness and generalizability of I2O visual localization methods. Experiments on the MGL dataset and CC validation benchmark have demonstrated that our OSMLoc achieves higher localization accuracy, while extensive experiments on the KITTI dataset highlight its application value in the autonomous driving area. Sequential localization experiments underscore the potential practical value. We hope that our OSMLoc will benefit the relevant communities in the field.
CRediT authorship contribution statement
Youqi Liao: Writing - original draft, Methodology, Data curation. Xieyuanli Chen: Writing – review & editing, Methodology, Conceptualization. Shuhao Kang: Writing - review & editing, Visualization. Jianping Li: Writing - review & editing, Supervision, Funding acquisition. Zhen Dong: Writing - review & editing, Supervision. Hongchao Fan: Writing - review & editing, Validation. Bisheng Yang: Supervision, Funding acquisition.
Acknowledgment
This study was jointly supported by the National Natural Science Foundation Project (No.42201477, No.42130105).